OpenStudy (anonymous):

15 over square root of 7 .

OpenStudy (anonymous):

15\[\sqrt{?}\]/7

OpenStudy (anonymous):

Okay, here's the setup\[15/\sqrt{7}\] Now we've looked at a couple of these. So, what do we need to multiply by to rationalize Lyssah?

OpenStudy (anonymous):

oops lol 15@7/7

OpenStudy (anonymous):

7 ?

OpenStudy (anonymous):

you have to rationalize the denom so there's no radicle. To do that you multiply the denom by itself, then the numerator. And that can't be simplified anymore. So almost anytime you have a problem with a radicle on the bottom, add it to the numerator and remove the radicle from the denominator. :)

OpenStudy (anonymous):

\[\sqrt{7}/\sqrt{7}\] that way we get ride of the sqrt on the bottom of the fraction, but at the same time we are just multiplying by something equal to 1 so it doesn't change the value, just how it looks